# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v2

@@ Appendix C Chain-of-Thought Keyword Categories
C.1. Example Chain-of-Thought Rationales
@@
Appendix D Supplementary Results
Figure 2 and Tables 10, 11 and 12 contain supporting tables that provide additional detail for the findings presented in the main text.
Table 10. Economic Scenario Green Transition Rates: Comparison of With Personas vs Without Personas. This table shows Green Transition selection rates in economic scenarios (where Tragedy is payoff-optimal). Each row shows a different model-variant combination. The table demonstrates that removing personas reduces Green Transition selection by 10-37 percentage points, indicating that personas bias models toward Green actions even in economic scenarios where Tragedy is optimal. This supports the finding that role identity bias prevents payoff-optimal equilibrium selection.
Model Variant With Persona Without Persona Difference
Llama-8B Hidden 45.0% (27/60) 26.7% (16/60) -18.3 pp
Mistral-7B Hidden 58.3% (35/60) 21.7% (13/60) -36.6 pp
Qwen-32B Hidden 31.7% (19/60) 15.0% (9/60) -16.7 pp
Qwen-7B Hidden 28.3% (17/60) 18.3% (11/60) -10.0 pp
Table 11. Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs. This table compares overall Nash equilibrium rates (across all scenarios and models) between hidden and visible payoffs conditions. Each row shows a different model-persona combination. The table reveals that with personas, hidden payoffs achieve equal or higher Nash rates than visible payoffs, suggesting reasoning scaffolding is more effective when role identity bias is present. Without personas, visible payoffs show large decreases for Mistral and Qwen models, but for Qwen, this is due to enabling Tragedy in economic scenarios (which is correct behavior).
Model Persona Hidden Visible Difference
Llama-8B Yes 77.4% (205/265) 77.0% (204/265) -0.4 pp
Llama-8B No 75.5% (200/265) 72.1% (191/265) -3.4 pp
Mistral-7B Yes 77.4% (205/265) 69.8% (185/265) -7.6 pp
Qwen-32B Yes 69.1% (183/265) 57.7% (153/265) -11.4 pp
Qwen-7B Yes 75.5% (200/265) 75.5% (200/265) 0.0 pp
Mistral-7B No 75.8% (201/265) 38.9% (103/265) -36.9 pp
Qwen-32B No 74.3% (197/265) 33.2% (88/265) -41.1 pp
Qwen-7B No 75.5% (200/265) 55.5% (147/265) -20.0 pp
Table 12. Persona Effect in Environmental Scenarios: Nash Equilibrium Rate Change (Without Personas vs With Personas). This table shows persona effects specifically for environmental scenarios. The structure is similar to Table 6(a), but focuses only on the 41 environmental scenarios. The table shows that removing personas generally reduces environmental Nash rates (negative values), with the largest effects for Mistral and Qwen models with visible payoffs (-47.8 $p$ and -58.0 $p$, respectively). This indicates that personas are particularly important for coordinating Green Transition in environmental scenarios, especially when explicit payoffs are provided.
Model Hidden Visible Average
Llama-8B -2.4 pp -6.3 pp -4.4 pp
Mistral-7B -2.0 pp -47.8 pp -24.9 pp
Qwen-32B +6.8 pp -58.0 pp -25.6 pp
Qwen 7B 0.0 pp -44.9 pp -22.5 pp
Appendix E Chain-of-Thought Analysis Figures
@@
E.1. Keyword Patterns Across Conditions
In figure 5, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning). This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed, and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
Figure 5. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model, detailing the distribution of keywords within the generated rationales.
E.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
@@
E.3. Model Comparison of Game-Theoretic Reasoning
All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition. However, this consistent high keyword usage does not translate to the ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas. This suggests that while keyword frequency is informative, the combination of strategic keywords and payoff visibility is necessary for payoff-optimal equilibrium selection, which only Qwen models achieve.
Figure 7. Game-Theoretic Keyword Frequency Across Models and Conditions. This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions.
E.4. Strategic Keywords - A Unique Pattern in Qwen 32B
Figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage. This unique pattern corresponds to Qwen 32B’s achievement of 90% Nash equilibrium in economic scenarios and 99.7% game-theoretic reasoning in rationales. Long-term keywords also appear more frequently for Qwen 32B (283 mentions) compared to other models in this condition. This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model architectures (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.

### v3

@@ I.1. Keyword Categories
I.2. Example Chain-of-Thought Rationales
@@
Appendix J Chain-of-Thought Analysis Figures
@@
J.1. Keyword Patterns Across Conditions
In figure 9, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning). This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
Figure 9. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model.
J.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
@@
J.3. Model Comparison of Game-Theoretic Reasoning
All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition. However, this consistent high keyword usage does not translate to ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas. This suggests that while keyword frequency is informative, the combination of strategic keywords and payoff visibility is necessary for payoff-optimal equilibrium selection, which only Qwen models achieve.
J.4. Strategic Keywords: Unique Pattern in Qwen 32B
The figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage. This unique pattern corresponds to Qwen 32B’s achievement of 90% Nash equilibrium in economic scenarios and 99.7% game-theoretic reasoning in rationales. Long-term keywords also appear more frequently for Qwen 32B (283 mentions) compared to other models in this condition. This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model families (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.